find 3 consecutive even integers such that five times the second equals twice the sum of the first and third
Answer by stanbon(48558)

find 3 consecutive even integers such that five times the second equals twice the sum of the first and third
-----------------------
1st: 2x
2nd: 2x+2
3rd: 2x+5
-----------------
Equation:
5(2x+2) = 2(2x + 2x+5)
10x + 10 = 2(5x+10)
10x + 10 = 10x + 20
0 = 10
That is a contradiction.
There are no consecutive even integers that meet the requirement.
================================================================
Cheers,
Stan H.